# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@ 3.4.1 Solving a second-order $q$-difference equation
$\Pi\upsilon=\exp\left(\sum_{m=0}^{Q-1}\Upsilon(q^{m}X)\right)=\exp\left(Q\varphi_{Q}(X)\right)\,,$ (3.111)
which is a direct consequence of Eq. (3.84a). Again, the sign ambiguity in Eq. (3.4.1) is resolved by the requirement in Eq. (3.109).
3.4.2 Solving a third-order $q$-difference equation
We will show how the matrix formalism of Section 3.4.1 can be applied to solve a third-order $q$-difference equation in closed form. In particular, let us return to the quantum $A$-polynomial equation satisfied by the resummed perturbative wave function $\chi^{(\alpha),\rm{WKB}}(u;\hbar)$ in Eq. (3.89) and assume it is of order three. As before, after appropriately accounting for the correction due to the exponential prefactor containing the functions $s^{(\alpha)}_{0,1}(x)$ in Eq. (3.87), the $q$-difference equation takes the form
$\Psi^{(\alpha),\rm{WKB}}(qX;q)+A(X)\Psi^{(\alpha),\rm{WKB}}(X;q)+B(X)\Psi^{(\alpha),\rm{WKB}}(q^{-1}X;q)+C(X)\Psi^{(\alpha),\rm{WKB}}(q^{-2}X;q)=0\,,$ (3.112)
where $A(X),B(X),C(X)$ are functions of $X$ and $q$. We can now introduce the function $\upsilon^{(\alpha)}(X)$ as in Eq. (3.91) and write Eq. (3.112) equivalently as
$\upsilon(X)\upsilon(q^{-1}X)\upsilon(q^{-2}X)+A(X)\upsilon(q^{-1}X)\upsilon(q^{-2}X)+B(X)\upsilon(q^{-2}X)+C(X)=0\,,$ (3.113)
where we have dropped the label $\alpha$ from the notation for simplicity. Let us further assume that $q$ is fixed as in Eq. (3.93) and define the the sequence of functions $\{\upsilon_{k}(X)\}$, $k\in{\mathbb{Z}}_{\geq 0}$, of $X$ and $q$ by
$\upsilon_{k}=\upsilon(q^{-k}X)\,.$ (3.114)
It follows from shifting $X$ into $q^{-k}X$ in Eq. (3.113) that they satisfy the system of equations
$\upsilon_{k}\upsilon_{k+1}\upsilon_{k+2}+A(q^{-k}X)\upsilon_{k+1}\upsilon_{k+2}+B(q^{-k}X)\upsilon_{k+2}+C(q^{-k}X)=0\,,\quad 0\leq k\leq Q-1\,.$ (3.115)
In order to solve the above system recursively, let us define the sequences $\{a^{(k)}(X)\}$, $\{b^{(k)}(X)\}$, and $\{c^{(k)}(X)\}$, $k\in{\mathbb{Z}}_{\geq 0}$, via the relations
@@ 3.4.2 Solving a third-order $q$-difference equation
$\upsilon_{0}\upsilon_{1}\cdots\upsilon_{k+1}+a^{(k)}(X)\upsilon_{k}\upsilon_{k+1}+b^{(k)}(X)\upsilon_{k+1}+c^{(k)}(X)=0\,,\quad k>1\,.$ (3.116d)
If we multiply Eq. (3.116d) by $\upsilon_{k+2}$ and apply Eq. (3.115), we obtain a recursion relation which can be written in matrix form as
$\begin{pmatrix}a^{(k+1)}(X)\\ b^{(k+1)}(X)\\ c^{(k+1)}(X)\end{pmatrix}=\begin{pmatrix}-A(q^{-k}X)&1&0\\ -B(q^{-k}X)&0&1\\ -C(q^{-k}X)&0&0\end{pmatrix}\begin{pmatrix}a^{(k)}(X)\\ b^{(k)}(X)\\ c^{(k)}(X)\end{pmatrix}\,,$ (3.117)
@@
${\cal M}_{Q}(X)=\prod_{k=1}^{Q}\begin{pmatrix}-A(q^{k}X)&1&0\\ -B(q^{k}X)&0&1\\ -C(q^{k}X)&0&0\end{pmatrix}=\begin{pmatrix}{\cal M}_{11}(X)&{\cal M}_{12}(X)&{\cal M}_{13}(X)\\ {\cal M}_{21}(X)&{\cal M}_{22}(X)&{\cal M}_{23}(X)\\ {\cal M}_{31}(X)&{\cal M}_{32}(X)&{\cal M}_{33}(X)\end{pmatrix}\,,$ (3.119)
@@
${\cal M}_{Q}(q^{-1}X)=\begin{pmatrix}-A(X)&1&0\\ -B(X)&0&1\\ -C(X)&0&0\end{pmatrix}{\cal M}_{Q}(X)\begin{pmatrix}-A(X)&1&0\\ -B(X)&0&1\\ -C(X)&0&0\end{pmatrix}^{-1}\,,$ (3.121)
and we obtain in this way
$\begin{pmatrix}-A{\cal M}_{12}+{\cal M}_{22}&-A{\cal M}_{13}+{\cal M}_{23}&\frac{A}{C}({\cal M}_{11}+A{\cal M}_{12}+B{\cal M}_{13})-\frac{1}{C}({\cal M}_{21}+A{\cal M}_{22}+B{\cal M}_{23})\\ -B{\cal M}_{12}+{\cal M}_{32}&-B{\cal M}_{13}+{\cal M}_{33}&\frac{B}{C}({\cal M}_{11}+A{\cal M}_{12}+B{\cal M}_{13})-\frac{1}{C}({\cal M}_{21}+A{\cal M}_{22}+B{\cal M}_{23})\\ -C{\cal M}_{12}&-C{\cal M}_{13}&({\cal M}_{11}+A{\cal M}_{12}+B{\cal M}_{13})\end{pmatrix}\,,$ (3.122)
where we have removed the explicit dependence on $X$ for compactness. Let us now take Eq. (3.116d) for $k=Q$, which gives
$\upsilon_{0}\upsilon_{1}(\Pi\upsilon+a^{(Q)}(X))+b^{(Q)}(X)\upsilon_{1}+c^{(Q)}(X)=0\,.$ (3.123)
Eq. (3.118) allows us to express the functions $a^{(Q)}(X)$, $b^{(Q)}(X)$, and $c^{(Q)}(X)$ in terms of the entries of the matrix $\mathcal{M}_{Q}(X)$, so that Eq. (3.123) becomes
@@
Shifting $X$ into $q^{-1}X$, using Eq. (3.115) with $k=0$, and applying the dictionary encoded in the matrix in Eq. (3.122), we find the second equation
@@
Subsequently shifting $X$ into $q^{-1}X$ in the equation above, again using Eqs. (3.115), (3.122), we obtain the third equation
@@
Eqs. (3.124), (3.125), (3.126) assemble into a cubic system which can be written in matrix form as
@@
$p(z)=\det(z\mathbb{I}_{3}-{\cal M}_{Q}^{\rm T})=z^{3}+d_{2}z^{2}+d_{1}z+d_{0}\,,$ (3.128)
where $\mathbb{I}_{3}$ is the 3x3 identity matrix and
@@
$\Pi\upsilon=-\frac{1}{3}\left(d_{2}+\Theta_{2}+\frac{\Theta_{0}}{\Theta_{2}}\right)\,,$ (3.130)
@@
$\Theta_{2}$ $={{\rm e}^{m{\pi{\rm i}\over 3}}\over 2}\sqrt[3]{\Theta_{1}+\sqrt{\Theta_{1}^{4}-4\Theta_{0}^{3}}}\,,\quad m=0,1,2\,.$ (3.131c)
The function $\upsilon_{0}$ is obtained, for example, by solving the system of Eqs. (3.125), (3.126), yielding
$\upsilon_{0}=\frac{{\cal M}_{22}{\cal M}_{33}-{\cal M}_{22}\Pi\upsilon-{\cal M}_{23}{\cal M}_{32}-{\cal M}_{33}\Pi\upsilon+(\Pi\upsilon)^{2}}{-{\cal M}_{12}{\cal M}_{33}+{\cal M}_{12}\Pi\upsilon+{\cal M}_{13}{\cal M}_{32}}\,,$ (3.132)
while the functions $\upsilon_{k}$, $k\in{\mathbb{Z}}_{>0}$, are computed by successively $q$-shifting $\upsilon_{0}$. This gives the solution to Eq. (3.113) for all rational values of $\hbar$. As we have done in Section 3.4.1, we can now fix the ambiguity in Eq. (3.130) and correspondingly in Eq. (3.132), that is, the value of $m=0,1,2$ in the definition of $\Theta_{2}$, by imposing the consistency of the solution with the original choice of a non-abelian classical branch labeled by $\alpha$. Specifically, we require that Eq. (3.109) holds in the case of $P=Q=1$ after accounting for the exponential prefactor in Eq. (3.87). Again, we conclude by observing that Eqs. (3.91), (3.111) yield
$\Upsilon(X)=\log\upsilon_{0}\,,\quad\varphi_{Q}(X)=\frac{1}{Q}\log\Pi\upsilon\,.$ (3.133)
@@ 4 Examples
In this section, we will illustrate how the conjectures and calculation methods of Section 3 apply to the two simplest hyperbolic knots, i.e., the figure-eight and the three-twist knots. We will first verify the integrality structure of Eq. (3.44) starting from the known decomposition into holomorphic and antiholomorphic blocks. We will then determine the wave function at rational values of $\hbar$ following the matrix formalism of Section 3.4 and cross-check the agreement of our results.
@@ 4.1 The figure-eight knot
The state integral of the figure-eight knot $\chi_{{\bf{4}}_{1}}(u;\hbar)$ can be expressed explicitly as bdp ; dimofte-levelk ; ggm2
$\chi_{{\bf{4}}_{1}}(u;\hbar)={\rm e}^{-\frac{\pi{\rm i}}{12}-\frac{2\pi{\rm i}}{3}c_{\mathsf{b}}^{2}}\Big{(}$ ${\rm e}^{\pi{\rm i}(-2u^{2}-4c_{\mathsf{b}}u)}{\cal J}({\rm e}^{4\pi{\mathsf{b}}u},{\rm e}^{2\pi{\mathsf{b}}u};q){\cal J}({\rm e}^{4\pi{\mathsf{b}}^{-1}u},{\rm e}^{2\pi{\mathsf{b}}^{-1}u};{\tilde{q}}^{-1})$ (4.134)
$+$ ${\rm e}^{\pi{\rm i}(u^{2}-2c_{\mathsf{b}}u)}{\cal J}({\rm e}^{2\pi{\mathsf{b}}u},{\rm e}^{-2\pi{\mathsf{b}}u};q){\cal J}({\rm e}^{2\pi{\mathsf{b}}^{-1}u},{\rm e}^{-2\pi{\mathsf{b}}^{-1}u};{\tilde{q}}^{-1})\Big{)}\,,$
@@
${\cal J}(x,y;q)=(yq;q)_{\infty}\sum_{n=0}^{\infty}(-1)^{n}\frac{q^{n(n+1)/2}x^{n}}{(q;q)_{n}(yq;q)_{n}}\,,$ (4.135)
while $(yq;q)_{\infty}$ and $(yq;q)_{n}$ denote the quantum dilogarithm and the $q$-shifted factorial, respectively. See Appendix C for their definitions. By means of Eq. (4.134), we can identify
$\phi^{(\rm{geom}),\rm{WKB}}(X;q)$ $=\log{\cal J}(X^{2},X;q)\,,$ (4.136a)
$\phi^{(\rm{conj}),\rm{WKB}}(X;q)$ $=\log{\cal J}(X,X^{-1};q)\,,$ (4.136b)
together with $C_{\rm{geom}}=C_{\rm{conj}}={\rm e}^{\frac{\pi{\rm i}}{4}}$ and
@@
Indeed, the two terms in the RHS of Eq. (4.134) represent the contributions coming from the geometric and conjugate branches, respectively. These are the only two non-abelian branches of the $\bm{4}_{1}$-knot, as described in Example 2.1. Moreover, we solve Eqs. (4.137a), (4.137b) order by order in $x$ and $\hbar$ and obtain that
$s^{(\rm{geom})}_{0}(x)$ $=\frac{\pi^{2}}{3}-\frac{x^{2}}{2}\,,\quad s^{(\rm{geom})}_{1}(x)=x\,,$ (4.138a)
$s^{(\rm{conj})}_{0}(x)$ $=\frac{\pi^{2}}{3}+\frac{x^{2}}{4}\,,\quad s^{(\rm{conj})}_{1}(x)=\frac{x}{2}\,.$ (4.138b)
In this way, we have fully determined the perturbative wave functions $\chi^{(\alpha),\rm{WKB}}(u;\hbar)$ of the figure-eight knot in the resummed form of Eq. (3.39). We verify successfully that they are annihilated by the $\hat{A}$-operator in Eq. (2.31), as expected.
Let us now test the conjectural integrality structure in Eq. (3.44) for both the geometric and the conjugate branches. Expanding the RHS of Eqs. (4.136a), (4.136b) gives
$\phi^{(\rm{geom}),\rm{WKB}}(X;q)$ $=\sum_{k\geq 1}\frac{q^{k}X^{k}}{k(q^{k}-1)}+\log\left(\sum_{n\geq 0}\frac{(-1)^{n}q^{n(n+1)/2}X^{2n}}{\prod_{i=1}^{n}(1-q^{i})(1-q^{i}X)}\right)\,,$ (4.139a)
$\phi^{(\rm{conj}),\rm{WKB}}(X;q)$ $=\sum_{k\geq 1}\frac{q^{k}X^{-k}}{k(q^{k}-1)}+\log\left(\sum_{n\geq 0}\frac{(-1)^{n}q^{n(n+1)/2}X^{2n}}{\prod_{i=1}^{n}(1-q^{i})(X-q^{i})}\right)\,.$ (4.139b)
Let us note that both terms on the RHS of Eq. (4.139a) for the geometric branch contribute with positive powers of $X$. For the conjugate branch, we note that the negative powers of $X$ arise entirely from the first term on the RHS of Eq. (4.139b), while a factor of $q^{\sigma_{{\rm conj},n}}$ with
$\sigma_{{\rm conj},n}=\left\lfloor\frac{n^{2}}{8}\right\rfloor\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (4.140)
appears in the denominator of the power series in $X$ obtained by expanding the second term. Let us stress that this negative power of $q$ arises from the factors $\prod_{i=1}^{n}(X-q^{i})^{-1}$ in the summand inside the logarithm, and it is absent from both the $X^{-1}$-series for the conjugate branch and the $X$-series for the geometric branch.
After extracting the coefficients $a^{(\alpha)}_{\pm n}(q)$, $n\in{\mathbb{Z}}_{>0}$, from the series for $\phi^{(\alpha),\rm{WKB}}(X;q)$, one can obtain the functions $D_{\pm s}^{(\alpha)}(q)$, $s\in{\mathbb{Z}}_{>0}$, by applying the formulae in Eqs. (3.46), (3.49). It is non-trivial that the resulting functions are polynomials with integer coefficients. In particular, the first few of them are
$D^{({\rm geom})}_{1}(q)$ $=D^{({\rm geom})}_{2}(q)=q\,,\quad D^{({\rm geom})}_{3}(q)=q^{2}\,,\quad D^{({\rm geom})}_{4}(q)=q^{3}\,,$ (4.141)
$D^{({\rm geom})}_{5}(q)$ $=q^{3}(1+q)\,,\quad D^{({\rm geom})}_{6}(q)=q^{4}(1+2q)\,,\quad D^{({\rm geom})}_{7}(q)=q^{4}(1+q)(1+q+q^{2})\,,$
@@
$D^{({\rm conj})}_{1}(q)$ $=0\,,\quad D^{({\rm conj})}_{2}(q)=D^{({\rm conj})}_{3}(q)=D^{({\rm conj})}_{4}(q)=-1\,,\quad D^{({\rm conj})}_{5}(q)$ $=-1-q\,,$ (4.142)
$D^{({\rm conj})}_{6}(q)$ $=-2-q\,,\quad D^{({\rm conj})}_{7}(q)=-(1+q)(1+q+q^{2})\,,$

### v2

@@ 3.4.1 Solving a second-order $\bm{q}$-difference equation
$\Pi\upsilon=\exp\left(\sum_{m=0}^{Q-1}\Upsilon(q^{m}X)\right)=\exp(Q\varphi_{Q}(X)),$ (3.73)
which is a direct consequence of equation (3.46a). Again, the sign ambiguity in equation (3.72) is resolved by the requirement in equation (3.71).
3.4.2 Solving a third-order $\bm{q}$-difference equation
We will show how the matrix formalism of Section 3.4.1 can be applied to solve a third-order $q$-difference equation in closed form. In particular, let us return to the quantum $A$-polynomial equation satisfied by the resummed perturbative wave function $\chi^{(\alpha),{\rm WKB}}(u;\hbar)$ in equation (3.51) and assume it is of order three. This is the case of the $\bm{5}_{2}$-knot, as shown in equation (2.36). As before, after appropriately accounting for the correction due to the exponential prefactor containing the functions $s^{(\alpha)}_{0,1}(x)$ in equation (3.49), the $q$-difference equation takes the form
$\Psi^{(\alpha),{\rm WKB}}(qX;q)+A(X)\Psi^{(\alpha),{\rm WKB}}(X;q)+B(X)\Psi^{(\alpha),{\rm WKB}}\bigl(q^{-1}X;q\bigr)$
$\qquad{}+C(X)\Psi^{(\alpha),{\rm WKB}}\bigl(q^{-2}X;q\bigr)=0,$ (3.74)
where $A(X)$, $B(X)$, $C(X)$ are functions of $X$ and $q$. We can now introduce the function $\upsilon^{(\alpha)}(X)$ as in equation (3.53) and write equation (3.74) equivalently as
$\upsilon(X)\upsilon\bigl(q^{-1}X\bigr)\upsilon\bigl(q^{-2}X\bigr)+A(X)\upsilon\bigl(q^{-1}X\bigr)\upsilon\bigl(q^{-2}X\bigr)+B(X)\upsilon\bigl(q^{-2}X\bigr)+C(X)=0,$ (3.75)
where we have dropped the label $\alpha$ from the notation for simplicity. Let us further assume that $q$ is fixed as in equation (3.55) and define the sequence of functions $\{\upsilon_{k}\}$, $k\in{\mathbb{Z}}_{\geq 0}$, of $X$ and $q$ by
$\upsilon_{k}=\upsilon\bigl(q^{-k}X\bigr).$ (3.76)
It follows from shifting $X$ into $q^{-k}X$ in equation (3.75) that they satisfy the system of equations
$\upsilon_{k}\upsilon_{k+1}\upsilon_{k+2}+A\bigl(q^{-k}X\bigr)\upsilon_{k+1}\upsilon_{k+2}+B\bigl(q^{-k}X\bigr)\upsilon_{k+2}+C\bigl(q^{-k}X\bigr)=0,$
$0\leq k\leq Q-1.$ (3.77)
In order to solve the above system recursively, let us define the sequences $\big\{a^{(k)}(X)\big\}$, $\big\{b^{(k)}(X)\big\}$, and $\big\{c^{(k)}(X)\big\}$, $k\in{\mathbb{Z}}_{\geq 0}$, via the relations
@@ 3.4.2 Solving a third-order $\bm{q}$-difference equation
$\upsilon_{0}\upsilon_{1}\cdots\upsilon_{k+1}+a^{(k)}(X)\upsilon_{k}\upsilon_{k+1}+b^{(k)}(X)\upsilon_{k+1}+c^{(k)}(X)=0,\qquad k>1.$ (3.78d)
If we multiply equation (3.78d) by $\upsilon_{k+2}$ and apply equation (3.77), we obtain a recursion relation which can be written in matrix form as
$\begin{pmatrix}a^{(k+1)}(X)\\ b^{(k+1)}(X)\\ c^{(k+1)}(X)\end{pmatrix}=\begin{pmatrix}-A\bigl(q^{-k}X\bigr)&1&0\\ -B\bigl(q^{-k}X\bigr)&0&1\\ -C\bigl(q^{-k}X\bigr)&0&0\end{pmatrix}\begin{pmatrix}a^{(k)}(X)\\ b^{(k)}(X)\\ c^{(k)}(X)\end{pmatrix}$, (3.79)
@@
${\cal M}_{Q}(X)=\prod_{k=1}^{Q}\begin{pmatrix}-A\bigl(q^{k}X\bigr)&1&0\\ -B\bigl(q^{k}X\bigr)&0&1\\ -C\bigl(q^{k}X\bigr)&0&0\end{pmatrix}=\begin{pmatrix}{\cal M}_{11}(X)&{\cal M}_{12}(X)&{\cal M}_{13}(X)\\ {\cal M}_{21}(X)&{\cal M}_{22}(X)&{\cal M}_{23}(X)\\ {\cal M}_{31}(X)&{\cal M}_{32}(X)&{\cal M}_{33}(X)\end{pmatrix}$, (3.81)
@@
${\cal M}_{Q}\bigl(q^{-1}X\bigr)=\begin{pmatrix}-A(X)&1&0\\ -B(X)&0&1\\ -C(X)&0&0\end{pmatrix}{\cal M}_{Q}(X)\begin{pmatrix}-A(X)&1&0\\ -B(X)&0&1\\ -C(X)&0&0\end{pmatrix}^{-1},$ (3.83)
which can be written as
$\left(\!\begin{matrix}-A{\cal M}_{12}\!+\!{\cal M}_{22}\!&\!-A{\cal M}_{13}\!+\!{\cal M}_{23}&\!{{A\over C}}({\cal M}_{11}\!+\!A{\cal M}_{12}\!+\!B{\cal M}_{13})\!-\!{{1\over C}}({\cal M}_{21}\!+\!A{\cal M}_{22}\!+\!B{\cal M}_{23})\vskip 2.84526pt\\ -B{\cal M}_{12}\!+\!{\cal M}_{32}\!&\!-B{\cal M}_{13}\!+\!{\cal M}_{33}&\!{{B\over C}}({\cal M}_{11}\!+\!A{\cal M}_{12}\!+\!B{\cal M}_{13})\!-\!{{1\over C}}({\cal M}_{21}\!+\!A{\cal M}_{22}\!+\!B{\cal M}_{23})\vskip 2.84526pt\\ -C{\cal M}_{12}&-C{\cal M}_{13}&({\cal M}_{11}\!+\!A{\cal M}_{12}\!+\!B{\cal M}_{13})\end{matrix}\!\right)\!,$ (3.84)
where we have removed the explicit dependence on $X$ for compactness. Let us now take equation (3.78d) for $k=Q$, which gives
$\upsilon_{0}\upsilon_{1}\bigl(\Pi\upsilon+a^{(Q)}(X)\bigr)+b^{(Q)}(X)\upsilon_{1}+c^{(Q)}(X)=0.$ (3.85)
Equation (3.80) allows us to express the functions $a^{(Q)}(X)$, $b^{(Q)}(X)$, and $c^{(Q)}(X)$ in terms of the entries of the matrix $\mathcal{M}_{Q}(X)$, so that equation (3.85) becomes
@@
Shifting $X$ into $q^{-1}X$, using equation (3.77) with $k=0$, and applying the dictionary encoded in the matrix in equation (3.84), we find the second equation
@@
Subsequently shifting $X$ into $q^{-1}X$ in the equation above, again using equations (3.77) and (3.84), we obtain the third equation
@@
Equations (3.86), (3.87), and (3.88) assemble into a cubic system which can be written in matrix form as
@@
$p(z)=\det\bigl(z\mathbb{I}_{3}-{\cal M}_{Q}^{\rm T}\bigr)=z^{3}+d_{2}z^{2}+d_{1}z+d_{0},$ (3.90)
where $\mathbb{I}_{3}$ is the $3\times 3$ identity matrix and
@@
$\Pi\upsilon=-{{1\over 3}}\left(d_{2}+\Theta_{2}+{{\Theta_{0}\over\Theta_{2}}}\right),$ (3.92)
@@
$\Theta_{2}$ $={{\rm e}^{m{\pi{\rm i}\over 3}}\over 2}\sqrt[3]{\Theta_{1}+\sqrt{\Theta_{1}^{4}-4\Theta_{0}^{3}}},\qquad m=0,1,2.$ (3.93c)
The function $\upsilon_{0}$ is obtained, for example, by solving the system of equations (3.87) and (3.88), yielding
$\upsilon_{0}={{{\cal M}_{22}{\cal M}_{33}-{\cal M}_{22}\Pi\upsilon-{\cal M}_{23}{\cal M}_{32}-{\cal M}_{33}\Pi\upsilon+(\Pi\upsilon)^{2}\over-{\cal M}_{12}{\cal M}_{33}+{\cal M}_{12}\Pi\upsilon+{\cal M}_{13}{\cal M}_{32}}},$ (3.94)
while the functions $\upsilon_{k}$, $k\in{\mathbb{Z}}_{>0}$, are computed by successively $q$-shifting $\upsilon_{0}$. This gives the solution to equation (3.75) for all rational values of $\hbar$. As we have done in Section 3.4.1, we can now fix the ambiguity in equation (3.92) and correspondingly in equation (3.94), that is, the value of $m=0,1,2$ in the definition of $\Theta_{2}$, by imposing the consistency of the solution with the original choice of a non-abelian classical branch labeled by $\alpha$. Specifically, we require that equation (3.71) holds in the case of $P=Q=1$ after accounting for the exponential prefactor in equation (3.49). Again, we conclude by observing that equations (3.53) and (3.73) imply
$\Upsilon(X)=\log\upsilon_{0},\qquad\varphi_{Q}(X)={{1\over Q}}\log\Pi\upsilon.$ (3.95)
@@ 4 Examples
In this section, we will illustrate how the conjectures and computational methods of Section 3 apply to the two simplest hyperbolic knots, i.e., the figure-eight ($\bm{4}_{1}$) and three-twist ($\bm{5}_{2}$) knots. We will first verify the integrality structure of equation (3.6) starting from the known decomposition into holomorphic and antiholomorphic blocks. We will then determine the wave function at rational values of $\hbar$ following the matrix formalism of Section 3.4 and cross-check the agreement of our results. Finally, we will include a short proof of admissibility for the examples at hand based on a theorem of Kontsevich and Soibelman [34].
@@ 4.1 The figure-eight knot
The state integral of the figure-eight knot $\chi_{{\bf{4}}_{1}}(u;\hbar)$ can be expressed as [5, 8, 17]
$\chi_{{\bf{4}}_{1}}(u;\hbar)={}$ ${\rm e}^{-{{\pi{\rm i}\over 12}}-{{2\pi{\rm i}\over 3}}c_{{\mathsf{b}}}^{2}}\bigl({\rm e}^{\pi{\rm i}(-2u^{2}-4c_{{\mathsf{b}}}u)}{\cal J}\bigl({\rm e}^{4\pi{\mathsf{b}}u},{\rm e}^{2\pi{\mathsf{b}}u};q\bigr){\cal J}\bigl({\rm e}^{4\pi{\mathsf{b}}^{-1}u},{\rm e}^{2\pi{\mathsf{b}}^{-1}u};{\tilde{q}}^{-1}\bigr)$
$+{\rm e}^{\pi{\rm i}(u^{2}-2c_{{\mathsf{b}}}u)}{\cal J}\bigl({\rm e}^{2\pi{\mathsf{b}}u},{\rm e}^{-2\pi{\mathsf{b}}u};q\bigr){\cal J}\bigl({\rm e}^{2\pi{\mathsf{b}}^{-1}u},{\rm e}^{-2\pi{\mathsf{b}}^{-1}u};{\tilde{q}}^{-1}\bigr)\bigr),$ (4.1)
@@
${\cal J}(x,y;q)=(yq;q)_{\infty}\sum_{n=0}^{\infty}(-1)^{n}{{q^{n(n+1)/2}x^{n}\over(q;q)_{n}(yq;q)_{n}}},$ (4.2)
while $(yq;q)_{\infty}$ and $(yq;q)_{n}$ denote the quantum dilogarithm and the $q$-shifted factorial, respectively. See Appendix C for their definitions. By means of equation (4.1), we can identify
$\phi^{({\rm geom}),{\rm WKB}}(X;q)=\log{\cal J}\bigl(X^{2},X;q\bigr),$ (4.3a)
$\phi^{({\rm conj}),{\rm WKB}}(X;q)=\log{\cal J}\bigl(X,X^{-1};q\bigr),$ (4.3b)
together with $\mathsf{C}_{{\rm geom}}=\mathsf{C}_{{\rm conj}}={\rm e}^{{{\pi{\rm i}\over 4}}}$ and
@@
Indeed, the two terms in the right-hand side of equation (4.1) represent the contributions coming from the geometric and conjugate branches, respectively. These are the only two non-abelian branches of the $\bm{4}_{1}$-knot, as described in Example 2.1. Moreover, we solve equations (4.4a) and (4.4b) order by order in $x$ and $\hbar$ and obtain that
$s^{({\rm geom})}_{0}(x)={{\pi^{2}\over 3}}-{{x^{2}\over 2}},\qquad$ $s^{({\rm geom})}_{1}(x)=x,$ (4.5a)
$s^{({\rm conj})}_{0}(x)={{\pi^{2}\over 3}}+{{x^{2}\over 4}},\qquad$ $s^{({\rm conj})}_{1}(x)={{x\over 2}}.$ (4.5b)
In this way, we have fully determined the perturbative wave functions $\chi^{(\alpha),{\rm WKB}}(u;\hbar)$ of the figure-eight knot in the resummed form of equation (3.1). We verify successfully that they are annihilated by the $\hat{A}-$operator in equation (2.32), as expected.
Let us now test the conjectural integrality structure in equation (3.6) for both the geometric and the conjugate branches. Expanding the right-hand side of equations (4.3a) and (4.3b) gives
$\phi^{({\rm geom}),{\rm WKB}}(X;q)=\sum_{k\geq 1}{{q^{k}X^{k}\over k\bigl(q^{k}-1\bigr)}}+\log\Bigg(\sum_{n\geq 0}{{(-1)^{n}q^{n(n+1)/2}X^{2n}\over\prod_{i=1}^{n}(1-q^{i})(1-q^{i}X)}}\Bigg),$ (4.6a)
$\phi^{({\rm conj}),{\rm WKB}}(X;q)=\sum_{k\geq 1}{{q^{k}X^{-k}\over k\bigl(q^{k}-1\bigr)}}+\log\Bigg(\sum_{n\geq 0}{{(-1)^{n}q^{n(n+1)/2}X^{2n}\over\prod_{i=1}^{n}(1-q^{i})(X-q^{i})}}\Bigg).$ (4.6b)
Let us note that both terms on the right-hand side of equation (4.6a) for the geometric branch contribute with positive powers of $X$. For the conjugate branch, we note that the negative powers of $X$ arise entirely from the first term on the right-hand side of equation (4.6b), while a factor of $q^{\sigma_{{\rm conj},n}}$ with
$\sigma_{{\rm conj},n}=\left\lfloor{{n^{2}\over 8}}\right\rfloor,\qquad n\in{\mathbb{Z}}_{>0},$ (4.7)
appears in the denominator of the power series in $X$ obtained by expanding the second term. Let us stress that this negative power of $q$ arises from the factors $\prod_{i=1}^{n}\bigl(X-q^{i}\bigr)^{-1}$ in the summand inside the logarithm, and it is absent from both the $X^{-1}$-series for the conjugate branch and the $X$-series for the geometric branch.
After extracting the coefficients $a^{(\alpha)}_{\pm n}(q)$, $n\in{\mathbb{Z}}_{>0}$, from the series for $\phi^{(\alpha),{\rm WKB}}(X;q)$, one obtains the functions $D_{\pm s}^{(\alpha)}(q)$, $s\in{\mathbb{Z}}_{>0}$, by applying the formulae in equations (3.8) and (3.11). It is non-trivial that the resulting functions are polynomials with integer coefficients. In particular, the first few of them are
$D^{({\rm geom})}_{1}(q)=D^{({\rm geom})}_{2}(q)=q,\qquad D^{({\rm geom})}_{3}(q)=q^{2},\qquad D^{({\rm geom})}_{4}(q)=q^{3},$
$D^{({\rm geom})}_{5}(q)=q^{3}(1+q),\qquad D^{({\rm geom})}_{6}(q)=q^{4}(1+2q),$
$D^{({\rm geom})}_{7}(q)=q^{4}(1+q)\bigl(1+q+q^{2}\bigr),$ (4.8)
@@
$D^{({\rm conj})}_{1}(q)=0,\qquad D^{({\rm conj})}_{2}(q)=D^{({\rm conj})}_{3}(q)=D^{({\rm conj})}_{4}(q)=-1,\qquad D^{({\rm conj})}_{5}(q)=-1-q,$
$D^{({\rm conj})}_{6}(q)=-2-q,\qquad D^{({\rm conj})}_{7}(q)=-(1+q)\bigl(1+q+q^{2}\bigr),$ (4.9)